Please help me with this. Find the slope of the line passing through the pair of points. (7, -3), (7,7)
answer: undefined slope

Answer by mtjmac(1)

using the formula for 2 point slope
y-y1 =(y2-y1/x2-x1)x-x1
y-(-3)= {7-(-3)/7-7}x-7
y+3=10/0 x-7
answer: undefined slope